# ReFoRCE: A Text-to-SQL Agent with Self-Refinement, Format Restriction, and Column Exploration

### v1

@@ Abstract
Text-to-SQL systems have unlocked easier access to critical data insights by enabling natural language queries over structured databases. However, deploying such systems in enterprise environments remains challenging due to factors such as large, complex schemas ($>3000$ columns), diverse SQL dialects (e.g., BigQuery, Snowflake) and sophisticated query requirements (e.g., transformation, analytics). Current state-of-the-art performance on the Spider 2.0 dataset — a benchmark built to mimic such complex environments — remains limited at 20%. Key limitations include inadequate instruction-following, poor long-context comprehension, weak self-refinement, and insufficient dialect-specific knowledge. To address these gaps, we propose ReFoRCE (Self-Refinement Agent with Format Restriction and Column Exploration) which introduces (1) table compression to mitigate long-context limitations (2) format restriction to ensure accurate answer format, and (3) iterative column exploration for enhanced schema understanding. Additionally, it employs self-refinement pipeline consisting of (1) parallelized workflows with voting mechanisms and (2) a Common Table Expression (CTE) based refinement approach to handle unresolved cases. ReFoRCE achieves state-of-the-art results scoring 26.69 on the Spider 2.0-Snow and scoring 24.50 on the Spider 2.0-Lite tasks.
@@ 1 Introduction
We evaluate our methods on two tasks of Spider 2.0 (Lei et al., 2024): Spider 2.0-Snow and Spider 2.0-Lite. Spider 2.0-Snow is unanimously in Snowflake SQL, while Spider 2.0-Lite includes examples in BigQuery and SQLite, additionally. ReFoRCE supports multiple dialects with minimal changes in prompts, making it versatile across various database systems. Our ReFoRCE agent achieves state-of-the-art results, scoring 26.69 on Spider 2.0-Snow and 24.50 on Spider 2.0-Lite, outperforming Spider-Agent’s score around 20. This demonstrates the effectiveness of our method in handling multiple dialects, nested columns and complex data types in Spider 2.0.
@@ 4.1 Experimental Setup
We conduct our experiments using the GPT family models, specifically GPT-4o (Achiam et al., 2023) and o1-preview (OpenAI, 2024). GPT-4o is used for testing some baselines, while o1-preview is employed for our method and Spider-Agent. We opted for o1-preview instead of the formal o1 version because o1-preview achieves better results for this particular task.
@@ 4.2 Evaluation Results
ReFoRCE (ours) o1-preview 26.69 24.50
Spider-Agent o1-preview 20.29 20.66
Spider-Agent GPT-4o 10.79 10.87
@@
The results in Table 1 highlight the superior performance of our method on the Spider 2.0-Snow and Spider 2.0-Lite datasets. Using the o1-preview model, our method achieves execution accuracy (EX) scores of 26.69 on Spider 2.0-Snow and 24.50 on Spider 2.0-Lite, significantly outperforming all other methods.
@@
Compared to Spider-Agent, which scores around 20 on both datasets using the same o1-preview model, our method achieves higher scores, highlighting its superior capability in solving complex cases in Text-to-SQL tasks. Additionally, methods using GPT-4o, including Spider-Agent, DAIL-SQL, and CHESS, perform poorly, with scores ranging from 0.00 to 5.68. This underscores the importance of our specialized approach, which is specifically designed to address the intricacies of Text-to-SQL tasks across diverse SQL dialects.
@@ 5 Conclusion
In this paper, we proposed ReFoRCE, a self-refinement framework for addressing the challenges of real-world Text-to-SQL tasks as exemplified by the Spider 2.0 dataset. By introducing techniques such as table information compression, format restriction, iterative column exploration, and CTE-based refinement, ReFoRCE effectively handles multiple SQL dialects, nested columns, and complex data types. Our approach achieved state-of-the-art performance, with scores of 26.69 on Spider 2.0-Snow and 24.50 on Spider 2.0-Lite, significantly outperforming the previous best results by Spider-Agent. These findings highlight the robustness and reliability of ReFoRCE in tackling real-world complexities.

### v2

@@ Abstract
Text-to-SQL systems have unlocked easier access to critical data insights by enabling natural language queries over structured databases. However, deploying such systems in enterprise environments remains challenging due to factors such as large, complex schemas ($>3000$ columns), diverse SQL dialects (e.g., BigQuery, Snowflake) and sophisticated query requirements (e.g., transformation, analytics). Current state-of-the-art performance on the Spider 2.0 dataset — a benchmark built to mimic such complex environments — remains limited at 20%. Key limitations include inadequate instruction-following, poor long-context comprehension, weak self-refinement, and insufficient dialect-specific knowledge. To address these gaps, we propose ReFoRCE (Self-Refinement Agent with Format Restriction and Column Exploration) which introduces (1) table compression to mitigate long-context limitations (2) format restriction to ensure accurate answer format, and (3) iterative column exploration for enhanced schema understanding. Additionally, it employs self-refinement pipeline consisting of (1) parallelized workflows with voting mechanisms and (2) a Common Table Expression (CTE) based refinement approach to handle unresolved cases. ReFoRCE achieves state-of-the-art results scoring 31.26 on the Spider 2.0-Snow and scoring 30.35 on the Spider 2.0-Lite tasks.
@@ 1 Introduction
We evaluate our methods on two tasks of Spider 2.0 (Lei et al., 2024): Spider 2.0-Snow and Spider 2.0-Lite. Spider 2.0-Snow is unanimously in Snowflake SQL, while Spider 2.0-Lite includes examples in BigQuery and SQLite, additionally. ReFoRCE supports multiple dialects with minimal changes in prompts, making it versatile across various database systems. Our ReFoRCE agent achieves state-of-the-art results, scoring 31.26 on Spider 2.0-Snow and 30.35 on Spider 2.0-Lite, outperforming Spider-Agent’s score around 20. This demonstrates the effectiveness of our method in handling multiple dialects, nested columns and complex data types in Spider 2.0.
@@ 4.1 Experimental Setup
We conduct our experiments using the GPT family models, specifically GPT-4o (Achiam et al., 2023) and o1-preview (OpenAI, 2024). GPT-4o is used for testing some baselines, while o1-preview is employed for our method and Spider-Agent. We opted for o1-preview instead of the formal o1 or o3-mini because o1-preview achieves better results for this particular task.
@@ 4.2 Evaluation Results
ReFoRCE (ours) o1-preview 31.26 30.35
Spider-Agent o1-preview 23.58 23.03
Spider-Agent GPT-4o 12.98 13.16
@@
The results in Table 1 highlight the superior performance of our method on the Spider 2.0-Snow and Spider 2.0-Lite datasets. Using the o1-preview model, our method achieves execution accuracy (EX) scores of 31.26 on Spider 2.0-Snow and 30.35 on Spider 2.0-Lite, significantly outperforming all other methods. [^1] Leaderboard data as of February 13, 2025.
@@
Compared to Spider-Agent, which scores around 23 on both datasets using the same o1-preview model, our method achieves higher scores, highlighting its superior capability in solving complex cases in Text-to-SQL tasks. Additionally, methods using GPT-4o, including Spider-Agent, DAIL-SQL, and CHESS, perform poorly, with scores ranging from 0.00 to 5.68. This underscores the importance of our specialized approach, which is specifically designed to address the intricacies of Text-to-SQL tasks across diverse SQL dialects.
@@ 5 Conclusion
In this paper, we proposed ReFoRCE, a self-refinement framework for addressing the challenges of real-world Text-to-SQL tasks as exemplified by the Spider 2.0 dataset. By introducing techniques such as table information compression, format restriction, iterative column exploration, and CTE-based refinement, ReFoRCE effectively handles multiple SQL dialects, nested columns, and complex data types. Our approach achieved state-of-the-art performance, with scores of 31.26 on Spider 2.0-Snow and 30.35 on Spider 2.0-Lite, significantly outperforming the previous best results by Spider-Agent. These findings highlight the robustness and reliability of ReFoRCE in tackling real-world complexities.